OpenStudy (anonymous):

Given: x + y = 6. If x={-15, -6, -1}, then which of the following sets of ordered pairs are solutions?

OpenStudy (hlambach):

Hey, my name is Hannah too! :) Anyways, what options of sets of ordered pairs were given?

OpenStudy (anonymous):

if 3+3=9 the i believe the answer is B

OpenStudy (anonymous):

{(-15, -21), (-6, -12), (-1, -7)} {(-15, 21), (-6, 12), (-1, 7)} {(-15, 21), (-6, 12), (-1, -7)} Theses are the answer choices.

OpenStudy (welshfella):

well hannah all you need do is substitute for the 3 values of x in the equation to find values of y

OpenStudy (hlambach):

Yeah, what welshfella said...

OpenStudy (welshfella):

for example put x = -15 into a + y = 6 -15 + y = 6 add 15 to both sides y = 6 +15 = 21 so we know that either b or c are the correct options as a has (-15,-21) so next try x = -6

OpenStudy (anonymous):

so is it b or c?

OpenStudy (welshfella):

yes - to find out which try plugging in the next value of x which is -6 and see what you get

OpenStudy (welshfella):

all you have to do is replace the x in x + y = 6 with -6 and work out what y is

OpenStudy (anonymous):

uhm... is still don't understand, but ok.

OpenStudy (welshfella):

-1 + y = 6 now solve for y

OpenStudy (anonymous):

y=7

OpenStudy (welshfella):

yeah = thats the third value - butthats ok now you should be able to choose the correct option b or c?

OpenStudy (welshfella):

find the one containing ( -1,7)

OpenStudy (anonymous):

Butthats? whats that even mean... and b

OpenStudy (welshfella):

lol but that's

OpenStudy (welshfella):

yes b is correct

OpenStudy (welshfella):

c has (-1,-7) which does not fit the calculations

OpenStudy (welshfella):

do you understand what we;ve done and why?

OpenStudy (anonymous):

yes...

OpenStudy (welshfella):

good

OpenStudy (anonymous):

fantastic

OpenStudy (anonymous):

i like your automobile

OpenStudy (anonymous):

well bye featheres thanks for helping me!!!

OpenStudy (anonymous):

b*tches

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

CHECK UR MESSAGES HO